# Mind the Gap: Removing the Discretization Gap in Differentiable Logic Gate Networks

### v1

@@ Abstract
Modern neural networks exhibit state-of-the-art performance on many existing benchmarks, but their high computational requirements and energy usage cause researchers to explore more efficient solutions for real-world deployment. Logic gate networks (LGNs) learns a large network of logic gates for efficient image classification. However, learning a network that can solve a simple problem like CIFAR-10 can take days to weeks to train. Even then, almost half of the network remains unused, causing a discretization gap. This discretization gap hinders real-world deployment of LGNs, as the performance drop between training and inference negatively impacts accuracy. We inject Gumbel noise with a straight-through estimator during training to significantly speed up training, improve neuron utilization, and decrease the discretization gap. We theoretically show that this results from implicit Hessian regularization, which improves the convergence properties of LGNs. We train networks 4.5×4.5$\times$4.5 × faster in wall-clock time, reduce the discretization gap by 98%, and reduce the number of unused gates by 100%.
@@ 1 Introduction
Figure 1: CIFAR-10 test accuracy comparison. Solid and dashed lines show discrete and soft performance, respectively. Gumbel LGNs (GLGN, red) demonstrate faster convergence and minimal discretization gap compared to Differentiable LGNs (DLGN, blue).
@@
Deep neural networks have reached human-level performance across a wide array of tasks. However, these advances come at the cost of immense computational demands during training and inference, limiting their deployment in many real-world environments. This has sparked growing interest in designing models that retain competitive accuracy while being more efficient.
@@
At their core, all computations on digital hardware reduce to Boolean operations: AND, OR, NOT, and so forth. This motivates: Can we express and execute machine learning models directly in the native language of hardware—namely, logic gates?
@@
One approach is Logic Gate Networks (LGNs), which replaces arithmetic computations with compositions of discrete logical operations, thereby enabling efficient inference. While inference with LGNs is efficient, training them poses significant challenges. To address this, Differentiable Logic Gate Networks (Differentiable LGNs) [1, 2] introduce continuous relaxations of both logical operations and gate selections, allowing the use of gradient-based optimization methods.
@@
Despite their potential, we identify and propose solutions to two major challenges. (1) Discretization gap: Differentiable LGNs rely on continuous relaxations during training, but final models must be discretized for inference. This mismatch may result in notable degradation, reducing model accuracy by $3\%$ on the same data. (2) Slow convergence: While LGNs are efficient at inference time, training remains slow due to the reliance on differentiable relaxations, making Differentiable LGNs converge substantially slower than standard neural networks.
@@
These challenges are interrelated. The gap arises because the final parameters, after training, must be discretized. Small parameter perturbations can significantly change performance if the loss landscape is sharp. A sharp loss landscape can also cause poor gradient signals, which impact the convergence speed, causing training to take much longer, while a smooth loss landscape can reduce the discretization gap and speed up convergence [3, 4]. Our central hypothesis is that smoother loss landscapes make LGN models more robust to discretization and facilitate faster and more stable training. Since the loss landscape is smoother, the gradient signal is better, thus making the networks converge faster. At the same time, the improved gradient signal also causes more neurons to collapse, thus reducing the impact of discretization.
@@
Figure 2: Overview figure. (a) Differentiable LGNs: Leftmost shows the internal structure of a node. We parameterize each node by weighing the 16 possible logic gates (shown below) and summing their output. This results in a hard, brittle, loss landscape, slowing convergence and increasing the discretization gap. (b) Gumbel LGNs: During training, we inject Gumbel noise on the 16 gate weights and select the logic gate with the highest weight. This results in a smoother loss landscape and aligns the training with the network at inference. This results in faster training and reduces the discretization gap. (c) Structure of Differentiable LGNs and Gumbel LGNs. Each neuron receives two inputs. The nodes in the final layer are aggregated by summation, thus producing class likelihoods. (d) Results of using Gumbel LGNs instead of Differentiable LGNs. We achieve up to 4.5×4.5$\times$4.5 × faster convergence (in wall-clock time), a 98% reduction in discretization gap, and 100% reduction of unused neurons.
@@
We propose Gumbel Logic Gate Networks (Gumbel LGNs), which inject Gumbel noise into the gate selection process using the Gumbel-Softmax trick. By injecting Gumbel noise into the gate selection process, we introduce stochasticity that flattens the loss landscape, improving the optimization process and reducing discretization sensitivity. Empirically, we find that Gumbel LGNs exhibit faster convergence and smaller discretization gaps than baseline Differentiable LGNs.
@@
To our knowledge, no prior work analyzes the discretization gap in LGNs at this scale, and draw formal connections to the smoothness of the loss landscape. We are also the first work that examines the discretization gap in parameter spaces similar to deep neural networks. While prior NAS studies explored search spaces up to size $10^{18}$, our experiments operate over vastly larger parameter spaces, exceeding $10^{3{,}600{,}000}$, and show that NAS techniques scale to this setting.
@@
  • Theoretical analysis. We prove that injecting Gumbel noise into Differentiable LGNs smooths their loss landscape by regularizing the Hessian’s trace, thereby reducing the discretization gap and accelerating convergence.
@@ 2 Background
Logic Gate Networks (LGNs) represent an entire network as a composition of discrete logic operations. In a learned network, each neuron in a hidden layer takes as input the value of two neurons (with output $a$ and $b$) in the previous layer [^1] At initialization, each neuron randomly picks which two neurons in the previous layer it uses for its inputs. and applies a fixed logic gate $h_{i}(a,b)$ to get its output. The final layer neurons are partitioned into $k$ disjoint groups $G_{i}$. Following Petersen et al. [1], we implement GroupSum, which computes class scores as $s_{i}=\frac{1}{\tau^{\mathrm{GS}}}\sum_{j\in G_{i}}a_{j}$, where $a_{j}$ is the binary activation of neuron $j$ and $\tau^{\mathrm{GS}}$ is a GroupSum temperature parameter. The class with the highest neuron activation count is predicted.
From Combinatorial Search to Differentiable Training
Directly searching for the best discrete gate assignments is infeasible due to the size of the search space, so Differentiable Logic Gate Networks (Differentiable LGNs) [1, 2] introduce a continuous relaxation. Each of the 16 possible binary gates $h_{i}(a,b)$ is replaced by a continuous surrogate (e.g. AND$(a,b)\mapsto a\cdot b$). In addition, each neuron maintains logits $\mathbf{z}\in\mathbb{R}^{16}$, which are initialized using a Gaussian, $\mathbf{z}\sim\mathcal{N}(0,1)^{16}$. After a softmax, these logits define a probability distribution over gates, and the neuron’s output is a weighted sum of the 16 gates:
@@
Ensuring each logic gate maps from a continuous domain $f:\left[0,1\right]^{2}\to\left[0,1\right]$. This “soft” network can be trained end-to-end with gradient descent.
@@
After training, Differentiable LGNs are discretized to LGNs by selecting the logic gate with the highest logit value, i.e., it uses $h_{i},i=\arg\max_{i}z_{i}$. We denote Differentiable LGNs evaluated in the differentiable setting (using Equation 1) as soft and otherwise as discrete.
@@
The Gumbel-Softmax trick offers an efficient and effective way to draw samples from a categorical distribution with class probabilities $\pi\in\Delta^{k}$ [5, 6, 7, 8]. Let $g\sim\mathrm{Gumbel}(0,1)$ distribution if $u\sim U(0,1)$ and $g=-\log(-\log u)$. We can then draw a sample $z$ from $\pi$ as the value for index $i$ given by Equation 2.
@@
We can make the argmax operation continuous and differentiable with respect to the class probabilities $\pi_{i}$, and generate $k$-dimensional sample vectors $y\in\mathbb{R}^{k}$ using a softmax with temperature $\tau$ as below:
@@ 3 Related Work
Of particular interest in this context are Differentiable Logic Gate Networks (LGNs), which have recently demonstrated state-of-the-art performance in image classification tasks [1, 2], as well as in rule extraction from observed cellular automata dynamics [25]. The convolutional variant’s code is not public; we therefore focus on the original LGN Petersen et al. [1]. Our proposed improvements target convergence behavior and are orthogonal to the architectural innovations of the convolutional LGN variant; hence, we expect them to be transferable without loss of generality. We refrain from comparisons to other efficient neural models, as such benchmarks were already comprehensively addressed in the works mentioned above by Petersen et al. [1, 2].
@@
A seminal contribution in this domain is Differentiable Architecture Search (DARTS) by Liu et al. [33], which introduces a softmax-based relaxation over discrete architectural choices, allowing end-to-end optimization through gradient descent. This principle strongly resonates with the soft gate selection mechanism employed in LGNs.
More recently, Chen and Hsieh [4] reduced the discretization gap of DARTS by introducing Smooth DARTS, which uses weight perturbations through uniform noise or adversarial optimization. These were shown to bias the optimization toward solutions with flatter minima and lower Hessian norm. This technique—often referred to as curvature regularization—reduces sensitivity to sharp local optima and enhances generalization.
@@
The works on LGNs by Petersen et al. [1, 2] do not explicitly draw connections to NAS, but the conceptual similarity is high. Both LGNs and DARTS use softmax-based weighting to choose between multiple candidate functions in a differentiable manner. A key distinction lies in the scale of the search space. Conventional NAS approaches typically explore search space sizes up to $10^{18}$ [34, 35, 36],while LGNs operate over exponentially larger spaces—$16^{6\cdot 64,000}\approx 10^{462,382}$ for MNIST and $\approx 10^{3,699,056}$ for CIFAR-10. This scale is enabled by the simplicity of logic operations, which have no learnable parameters. Thus, LGNs demonstrate the viability of differentiable NAS at previously unexplored scales.
@@
A parallel line of research focuses on improving generalization by minimizing the sharpness of the loss landscape. Motivated by prior theoretical works on generalization and flat minima [38, 39, 40], Foret et al. [3] introduced Sharpness-Aware Minimization (SAM) in Equation 4. This technique explicitly seeks flat minima by optimizing the worst-case loss within a perturbation neighborhood.
@@
where $L_{\mathcal{S}}(\bm{w})$ is a loss function over a training set $\mathcal{S}$ of training samples evaluated for model parameters $\bm{w}$. $p\in[1,\infty[$ is the $p$-norm used (usually $p=2$) and $\rho>0$ is a hyperparameter [3]. Since its introduction, SAM has inspired numerous follow-up studies focused on improving computational efficiency [41, 42, 43, 44, 45, 46] as well as providing theoretical insights into its efficacy [47, 48, 49, 50]. We refer to Appendix D for a detailed description of SAM.
@@ 4 Gumbel Logic Gate Networks
We introduce Gumbel Logic Gate Networks (Gumbel LGNs), which employ discrete sampling of logic gates via the Gumbel-Softmax trick with a straight-through (ST) estimator [7, 8]. While conventional Differentiable LGNs maintain a convex combination of gates throughout training and prune to hard selections at inference time, Gumbel LGNs resemble inference-time behavior directly during training by stochastically selecting individual gates per forward pass. We perturb the gate logits with Gumbel noise and select the most probable gate, following the argmax operation (cf. Equation 2). During backpropagation, the non-differentiable argmax is approximated using the Gumbel-Softmax (Equation 3), enabling end-to-end training.
This approach is motivated by two key observations: (1) Implicit smoothening via noise: Injecting Gumbel noise during the forward pass introduces a form of stochastic smoothing, effectively averaging over local perturbations of the loss surface. As we show, this process approximates a curvature-penalizing loss that favors flatter minima and smaller Hessian norm. In addition, this is known to correlate with improved generalization [3, 4]. (2) Inference-time alignment: In Differentiable LGNs, the training objective is misaligned with inference behavior, as training relies on weighted combinations of gates that are ultimately discarded. This discrepancy harms generalization. In contrast, Gumbel LGNs train under the same discrete selection mechanism, which is used at inference.
@@
As with Differentiable LGNs, we model each neuron as a distribution over binary, relaxed logic gates $\mathcal{S}=\{h_{1},h_{2},\dots,h_{16}\}$, where each gate $h_{i}:[0,1]^{2}\rightarrow[0,1]$ operates on relaxed Boolean inputs. We associate each gate $i$ with logit $z_{i}$, and the gate has weight $\pi_{i}^{Gumbel}$ from Equation 3. The output of the neuron with inputs $(a,b)$ is then:
@@
$\tau>0$ is a temperature parameter controlling the sharpness of the distribution. As $\tau\to 0$, the distribution increasingly peaks around the maximum-logit index [7]. This relaxation enables end-to-end differentiability while encouraging the network to commit to discrete logic gates during training.
We employ a straight-through (ST) estimator to bridge the discretization gap between the continuous relaxation used during training and the hard decisions required during inference. In this formulation, each neuron selects a single logic gate in the forward pass via a hard (non-differentiable) choice, while gradients are estimated through a soft relaxation in the backward pass. See Appendix E for pseudo-code implementation of the training process.
Concretely, during the forward pass, we sample Gumbel noise $\mathbf{g}\sim\mathrm{Gumbel}(0,1)^{16}$ and compute:
@@
using the gate $h_{k}$ with maximum perturbed logit. During the backward pass, we use the soft Gumbel-Softmax relaxation (Equation 5) to compute gradients, effectively treating the hard output as if it were differentiable:
@@
This ST estimator mechanism encourages the network to make discrete decisions and allows end-to-end optimization via backpropagation. See Figure 2 or Figure 9 in Appendix A for visualizations.
@@
Lemma 1 (Gumbel-Smoothing).
@@
See Appendix B. ∎
Intuitively, by injecting Gumbel noise during training, we encourage the optimizer to find parameters that are robust to small perturbations. This results in flatter loss landscapes, and reduces the sensitivity to parameter discretization when switching to inference mode. The expected loss scales as
@@
So as the temperature $\tau$ increases, the coefficient $1/\tau^{2}$ decreases, reducing the degree of implicit smoothing. We illustrate this with two representative choices:
@@
Figure 3: Performance of Gumbel LGNs and Differentiable LGNs on CIFAR-10. Left: Test accuracy for the default network with 12 layers and a width of $256$k. Right: Discretization gap for various depths. Differentiable LGNs experience larger gaps and slower reduction as the depth increases. In contrast, Gumbel LGNs have consistently low gaps and fast reduction as the network depth increases.

### v2

@@ Abstract
Modern neural networks exhibit state-of-the-art performance on many existing benchmarks, but their high computational requirements and energy usage cause researchers to explore more efficient solutions for real-world deployment. Differentiable logic gate networks (DLGNs) learns a large network of logic gates for efficient image classification. However, learning a network that can solve simple problems like CIFAR-10 or CIFAR-100 can take days to weeks to train. Even then, almost half of the neurons remains unused, causing a discretization gap. This discretization gap hinders real-world deployment of DLGNs, as the performance drop between training and inference negatively impacts accuracy. We inject Gumbel noise with a straight-through estimator during training to significantly speed up training, improve neuron utilization, and decrease the discretization gap. We theoretically show that this results from implicit Hessian regularization, which improves the convergence properties of DLGNs. We train networks $4.5\times$ faster in wall-clock time, reduce the discretization gap by 98%, and reduce the number of unused gates by 100%.
@@ 1 Introduction
Figure 1: CIFAR-10 test accuracy comparison. Solid and dashed lines show discrete and soft performance, respectively. Gumbel LGNs (GLGN, red) demonstrate faster convergence and minimal discretization gap compared to DLGNs (DLGN, blue).
@@
Deep neural networks have reached human-level performance across a wide array of tasks [1, 2]. However, these advances come at the cost of immense computational demands during training and inference, limiting their deployment in many real-world environments [3, 4, 5, 6]. This has sparked growing interest in designing models that retain competitive accuracy while being more efficient [7, 8, 9, 10, 11].
@@
At their core, all computations on digital hardware reduce to Boolean operations such as AND, OR, and NOT. This motivates the question: Can we express and execute machine learning models directly in the native language of hardware - namely, logic gates?
@@
One approach is logic gate networks (LGNs), which replaces arithmetic computations with compositions of discrete logical operations, thereby enabling efficient inference. While inference with LGNs is efficient, training them poses significant challenges. To address this, differentiable logic gate networks (DLGNs) [7, 9] introduce continuous relaxations of both logical operations and gate selections, allowing the use of gradient-based optimization methods.
@@
Despite their potential, we identify and propose solutions to two major challenges. (1) Discretization gap: DLGNs rely on continuous relaxations during training, but final models must be discretized for inference. This mismatch may result in notable degradation, reducing model accuracy by $3\%$ on the same data. (2) Slow convergence: While LGNs are efficient at inference time, training DLGNs remains slow due to the reliance on differentiable relaxations, making DLGNs converge substantially slower than standard neural networks.
@@
These challenges are interrelated. The gap arises because the final parameters, after training, must be discretized. Small parameter perturbations can significantly change performance if the loss landscape is sharp [12]. A sharp loss landscape can also cause poor gradient signals, which impact the convergence speed, causing training to take much longer, while a smooth loss landscape can reduce the discretization gap and speed up convergence [13, 14, 15, 16]. Our central hypothesis is that smoother loss landscapes make DLGN models more robust to discretization and facilitate faster and more stable training. Since the loss landscape is smoother, the gradient signal is better, thus making the networks converge faster. At the same time, the improved gradient signal also causes more neurons to collapse, thus reducing the impact of discretization.
@@
Figure 2: Overview figure. (a) DLGNs: Leftmost shows the internal structure of a node. We parameterize each node by weighing the 16 possible logic gates (shown below) and summing their output. This results in a hard, brittle, loss landscape, slowing convergence and increasing the discretization gap. (b) Gumbel LGNs: during training, we inject Gumbel noise on the 16 gate weights and select the logic gate with the highest weight. This results in a smoother loss landscape and aligns the training with the network at inference. This results in faster training and reduces the discretization gap. (c) Structure of DLGNs and Gumbel LGNs. Each neuron receives two inputs. The nodes in the final layer are aggregated by summation, thus producing class likelihoods. (d) Results of using Gumbel LGNs instead of DLGNs. We achieve up to $4.5\times$ faster convergence (in wall-clock time), a 98% reduction in discretization gap, and 100% reduction of unused neurons.
@@
We propose Gumbel Logic Gate Networks (Gumbel LGNs), which inject Gumbel noise into the gate selection process using the Gumbel-Softmax trick. By injecting Gumbel noise into the gate selection process, we introduce stochasticity that flattens the loss landscape, improving the optimization process and reducing discretization sensitivity. Empirically, we find that Gumbel LGNs exhibit faster convergence and smaller discretization gaps than baseline DLGNs.
@@
To our knowledge, no prior work analyzes the discretization gap in DLGNs, and draw formal connections to the smoothness of the loss landscape. NAS and DARTS typically studies explored search spaces up to size $10^{18}$, while our experiments operate over vastly larger parameter spaces, exceeding $10^{3{,}600{,}000}$, and show that NAS techniques scale to this setting.
@@
  • Theoretical analysis. We prove that injecting Gumbel noise into DLGNs smooths their loss landscape by regularizing the Hessian’s trace, thereby reducing the discretization gap and accelerating convergence.
@@ 2 Background
Logic Gate Networks (LGNs) represent an entire network as a composition of discrete logic operations, such as AND, OR, and XOR. LGNs consist of several layers, where each layer contains $n$ fixed logic gates, referred to as neurons, that work in parallel. $n$ is the width of the layer. Each neuron in a layer takes as input the output of two (random) neurons in the previous layer. We use the GroupSum operation to get class scores as done by Petersen et al. [7]. For classifying the input into $k$ classes, the neurons in the final layer are grouped into disjoint groups with one group per class. The outputs of the neurons in group $i$ are summed to give the likelihood of class $i$. Mathematically this is expressed as $s_{i}=\frac{1}{\tau^{\mathrm{GS}}}\sum_{j\in G_{i}}a_{j}$, where $G_{i}$ is the $i$’th group, $a_{j}$ is the output of neuron $j$, and $\tau^{\mathrm{GS}}$ is a temperature parameter. We then predict the class with the highest score. For more details, see [17].
Differentiable Training
Directly searching for the best discrete gate assignments is infeasible due to the size of the search space, so Differentiable Logic Gate Networks (DLGNs) [7, 9] introduce a continuous relaxation. Each of the 16 possible binary gates $h_{i}(a,b)$ is replaced by a continuous surrogate (e.g. AND(a,b)ab($\mapsto a\cdot b)$. In addition, each neuron maintains logits $\mathbf{z}\in\mathbb{R}^{16}$, which are initialized using a Gaussian, $\mathbf{z}\sim\mathcal{N}(0,1)^{16}$. After a softmax, these logits define a probability distribution over gates, and the neuron’s output is a weighted sum of the 16 gates:
@@
Ensuring each logic gate maps from a continuous domain $f:\left[0,1\right]^{2}\to\left[0,1\right]$. This “soft” network can be trained end‐to‐end with gradient descent.
@@
After training, DLGNs are discretized to LGNs by selecting the logic gate with the highest logit value, i.e., it uses $h_{i},i=\arg\max_{i}z_{i}$. We denote DLGNs evaluated in the differentiable setting (using Equation˜1) as soft and otherwise as discrete.
@@
The Gumbel-Softmax trick offers an efficient and effective way to draw samples from a categorical distribution with class probabilities $\pi\in\Delta^{k}$ [18, 19, 20, 21]. Let $g\sim\mathrm{Gumbel}(0,1)$ distribution if $u\sim U(0,1)$ and $g=-\log(-\log u)$. We can then draw a sample $z$ from $\pi$ as the value for index $i$ given by Equation˜2.
@@
We can make the argmax operation continuous and differentiable with respect to the class probabilities $\pi_{i}$, and generate $k$-dimensional sample vectors $y\in\mathbb{R}^{k}$ using the Softmax with temperature $\tau$ as below:
@@ 3 Related Work
Of particular interest in this context are differentiable logic gate networks (DLGNs), which have demonstrated state-of-the-art performance in image classification tasks [7, 9], as well as in rule extraction from observed cellular automata dynamics [17]. Our proposed improvements target convergence behavior and are orthogonal to the architectural innovations of the convolutional DLGN variant [9]; hence, we expect them to be transferable without loss of generality. We refrain from comparisons to other efficient neural models, as such benchmarks were already comprehensively addressed in the works mentioned above by Petersen et al. [7, 9].
Kim [38] extend DLGNs by modifying the neuron parameterization introduced in Petersen et al. [7]. Specifically, they normalize the logits and remove the temperature parameter $\tau$. In addition, they employ a straight-through (ST) estimator [39] and injectGumbel noise with a learnable scale parameter. However, Kim [38] do not analyze the discretization gap or study the convergence behavior of DLGNs. Moreover, they draw no formal connection between Gumbel noise and its role as an implicit Hessian regularizer or loss smoothener.
@@
A seminal contribution in this domain is Differentiable Architecture Search (DARTS) by Liu et al. [47], which introduces a softmax-based relaxation over discrete architectural choices, allowing end-to-end optimization through gradient descent. This principle strongly resonates with the soft gate selection mechanism employed in DLGNs.
More recently, Chen and Hsieh [16] reduced the discretization gap of DARTS by introducing Smooth DARTS, which uses weight perturbations through uniform noise or adversarial optimization. These were shown to bias the optimization toward solutions with flatter minima and lower Hessian norm. This technique often referred to as curvature regularization reduces sensitivity to sharp local optima and enhances generalization.
@@
[…]
@@
A parallel line of research focuses on improving generalization by minimizing the sharpness of the loss landscape. Motivated by prior theoretical works on generalization and flat minima [14, 52, 53], Foret et al. [15] introduced Sharpness-Aware Minimization (SAM) in Equation˜4. This technique explicitly seeks flat minima by optimizing the worst-case loss within a perturbation neighborhood.
@@
where $L_{\mathcal{S}}(\bm{w})$ is a loss function over a training set $\mathcal{S}$ of training samples evaluated for model parameters $\bm{w}$. $p\in[1,\infty[$ is the $p$-norm used (usually $p=2$) and $\rho>0$ is a hyperparameter [15]. Since its introduction, SAM has inspired numerous follow-up studies focused on improving computational efficiency [54, 55, 56, 57, 58, 59] as well as providing theoretical insights into its efficacy [60, 61, 62, 63]. We refer to Appendix˜D for a detailed description of SAM.
@@ 4 Gumbel Logic Gate Networks
We introduce Gumbel Logic Gate Networks (Gumbel LGNs), which employ discrete sampling of logic gates via the Gumbel-Softmax trick [20, 21] with a straight-through (ST) estimator [39]. While conventional DLGNs maintain a convex combination of gates throughout training and prune to hard selections at inference time, Gumbel LGNs resemble inference-time behavior directly during training by stochastically selecting individual gates per forward pass. We perturb the gate logits with Gumbel noise and select the most probable gate, following the argmax operation (cf. Equation˜2). During backpropagation, the non-differentiable argmax is approximated using the Gumbel-Softmax (Equation˜3), enabling end-to-end training.
This approach is motivated by two observations: (1) Implicit smoothening via noise: injecting Gumbel noise during the forward pass introduces a form of stochastic smoothing, effectively averaging over local perturbations of the loss surface. As we show, this process approximates a curvature-penalizing loss that favors flatter minima and smaller Hessian norm. In addition, this is known to correlate with improved generalization [15, 16]. (2) Inference-time alignment: In DLGNs, the training objective is misaligned with inference behavior, as training relies on weighted combinations of gates that are ultimately discarded. This discrepancy harms generalization. In contrast, Gumbel LGNs train under the same discrete selection mechanism, which is used at inference.
@@
As with DLGNs, we model each neuron as a distribution over binary, relaxed logic gates $\mathcal{S}=\{h_{1},h_{2},\dots,h_{16}\}$, where each gate $h_{i}:[0,1]^{2}\rightarrow[0,1]$ operates on relaxed Boolean inputs. We associate each gate $i$ with logit $z_{i}$, and the gate has weight $\pi_{i}^{Gumbel}$ from Equation˜3. The output of the neuron with inputs $(a,b)$ is then given below.
@@
where $\tau>0$ is a temperature parameter controlling the sharpness of the distribution. As $\tau\to 0$, the distribution increasingly peaks around the maximum-logit index [20]. This relaxation enables end-to-end differentiability while encouraging the network to commit to discrete logic gates during training.
We employ a straight-through (ST) estimator to bridge the discretization gap between the continuous relaxation used during training and the hard decisions required during inference. In this formulation, each neuron selects a single logic gate in the forward pass via a hard (non-differentiable) choice, while gradients are estimated through a soft relaxation in the backward pass. See Appendix˜E for pseudo-code implementation of the training process. Concretely, during the forward pass, we sample Gumbel noise $\mathbf{g}\sim\mathrm{Gumbel}(0,1)^{16}$ and compute:
@@
using the gate $h_{k}$ with maximum perturbed logit. During the backward pass, we use the soft Gumbel-Softmax relaxation (Equation˜5) to compute gradients, effectively treating the hard output as if it were differentiable:
@@
This ST estimator mechanism encourages the network to make discrete decisions and allows end-to-end optimization via backpropagation. See Figure˜2 or Figure˜11 in Appendix˜A for visualizations.
@@
Lemma 1 (Gumbel‐Smoothing).
Let $\mathcal{L}:\mathbb{R}^{16}\to\ \mathbb{R}$ be twice continuously differentiable (with Lipschitz Hessian), and let $\mathbf{z}\in\mathbb{R}^{16},\mathbf{g}\sim\mathrm{Gumbel}(0,1)^{16}$. Consider
@@
and set $\mathbf{a}=\mathbf{z}/\tau$ and $f(\mathbf{a})=\mathcal{L}(\mathrm{softmax}(\mathbf{a}))$. We then get the expression below.
$J(\mathbf{z})=\mathcal{L}(\mathrm{softmax}(\mathbf{z}/\tau))+\frac{\pi^{2}}{12\tau^{2}}\mathrm{tr}(H_{f}(\mathbf{z}/\tau))+O\left(\tau^{-3}\right).$
@@
See Appendix˜B. ∎
Intuitively, by injecting Gumbel noise during training, we encourage the optimizer to find parameters that are robust to small perturbations. This results in flatter loss landscapes and reduces the sensitivity to parameter discretization when switching to inference mode [12]. The expected loss scales with
@@
As the temperature $\tau$ increases, the coefficient $1/\tau^{2}$ decreases, reducing the degree of implicit smoothing. We illustrate this with two representative choices: